OpenStudy (anonymous):

To change from degrees celsius to degrees fahrenheit, you can use the equation f=9/5c+32. If the temperature is at least 55 degrees fahrenheit, what inequality can you use to find the temperature in celsius? a) 55> 9/5c+32 b) f>9/5 x 55+32 c) 55<9/5c+32 d) f<9/5 x 55+32

zepdrix (zepdrix):

\[\Large\bf\sf f=\frac{9}{5}c+32\] If the temperature `f` is `at least` 55, this translates to `f` is `greater than or equal to` 55. \[\Large\bf\sf \color{royalblue}{f}\ge55\]Plugging in our function for f gives, \[\Large\bf\sf \color{royalblue}{\frac{9}{5}c+32}\ge55\]

zepdrix (zepdrix):

Err hmm.. I guess your options have the 55 written on the left side..

zepdrix (zepdrix):

So if we flip it around to read it from the other direction, \[\Large\bf\sf 55\le \frac{9}{5}c+32\] This look like any of them? :o

OpenStudy (anonymous):

Thank you! @zepdrix